# The immersed boundary method: an accelerated SIMPLE approach for moving bodies

### v1

@@ 4 Theory: a spectral bound on the preconditioned system
In this section, we justify taking the Laplacian as a preconditioner for the primal Schur complement (18) of the coupled pressure-forces system (14). A good preconditioner should resemble the original matrix, in the sense of spectral equivalence, i.e., the spectrum of the preconditioned system should be bounded. Theorem 4.1 shows this property.
@@
has only solutions that satisfy $\lambda\in[0,1]$. To bound $\lambda$, we bound the generalized Rayleigh quotient
$\frac{\left\langle B^{T}C^{-1}B{\bf v},{\bf v}\right\rangle}{\left\langle L{\bf v},{\bf v}\right\rangle}=\frac{\left\langle G^{T}PG{\bf v},{\bf v}\right\rangle}{\left\langle G^{T}G{\bf v},{\bf v}\right\rangle}=\frac{\left\langle PG{\bf v},G{\bf v}\right\rangle}{\left\langle G{\bf v},G{\bf v}\right\rangle},\quad{\bf v}\in\mathbb{R}^{m}$ (27)
where $P\coloneqq R(R^{T}R)^{-1}R^{T}$ is an orthogonal projector (namely, $P^{2}=P$ and $P^{T}=P$). Hence, the only eigenvalues of $P$ are 0 and 1, and $\text{Null}(P)=\text{Range}(P)^{\perp},$ see, e.g., [28]. Particularly, every vector can be uniquely decomposed into a sum of orthogonal vectors in the nullspace and in the range of $P$, e.g., we can write $G{\bf v}={\bf v}_{1}+{\bf v}_{2}$, where ${\bf v}_{1}\in\text{Null}(P)$, ${\bf v}_{2}\in\text{Range}(P)$ and $\left\langle{\bf v}_{1},{\bf v}_{2}\right\rangle=0.$ Consequently,
@@
$10\leq\frac{\left\langle PG{\bf v},G{\bf v}\right\rangle}{\left\langle G{\bf v},G{\bf v}\right\rangle}=\frac{\|{\bf v}_{2}\|^{2}}{\|{\bf v}_{1}+{\bf v}_{2}\|^{2}}=\left(\frac{\|P({\bf v}_{1}+{\bf v}_{2})\|}{\|{\bf v}_{1}+{\bf v}_{2}\|}\right)^{2}\leq\|P\|^{2}_{2}=\left(\rho(P)\right)^{2}=$ 1 (29)
@@
The assumption of symmetry is not particularly restrictive, as it is natural to take the same delta functions for the regularization and interpolation, and the same discretization for the gradient and divergence. In practice, some implementation of the possible BC might lead to a non-symmetric Laplacian. Nevertheless, even when the Laplacian slightly differs from $G^{T}G$, the spectral result of Theorem 4.1 seems to hold in practice, as demonstrated in the next section.
@@
As an immediate corollary of Theorem 4.1, the error iteration matrix $T=I_{m}-L^{-1}S_{p}$ of the preconditioned system satisfies $\rho(T)\leq 1$. When this inequality is strong, convergence is guaranteed. A Krylov iterative method typically further accelerates the convergence and might converge in just a few iterations even when the spectral radius of the error iteration matrix is even greater than 1. However, the convergence rate of a Krylov iterative method is determined by the scattering of the eigenvalues, which can depend on the specific configuration. In typical cases, $B$ is extremely rank deficient, because the number of Lagrangian points is very small compared to the number of Eulerian cells. In such a case,
@@ 5 Results
5.1 Verifying the preconditioning method by two-dimensional experiments
The aim of this subsection is to validate the theoretical result of Theorem 4.1 numerically, and to demonstrate certain condition number properties. For convenience in demonstrating these properties, we chose relatively small 2D cases. The numerical experiments were conducted using a model problem consisting of a stationary cylinder placed at the center of a lid-driven cavity. Conditioning analysis was performed on the matrices $L$, $S_{p}$ and $\tilde{S}_{p}$ obtained from the numerical simulations.
In the first experiment we demonstrate the spectral result of Theorem 4.1, reading $L^{-1}S_{p}\subseteq[1,2]$. Table 1, which presents the spectrum’s support $[\lambda_{min},\lambda_{max}]$, shows that $\text{spec}(L^{-1}S_{p})\subseteq[1,2]$, as predicted theoretically. Since the scalar approximation of $C$ is crucial for the efficiency of our method, we also measure the spectrum’s support for $L^{-1}\tilde{S}_{p}$, with $\tilde{S}_{p}$ from Eq. (21). In this case, most of the eigenvalues satisfy $\lambda\in[1,2]$, and although some eigenvalues slightly exceed 2, the maximal eigenvalue tends to 2 as the grid resolution increases. It can be explained by that fact that the quality of the scalar approximation $C\approx\frac{1}{2}I$ improves with the grid resolution, as shown in [16].
Spectrum’s support for the preconditioned system
Grid size (cells) $24\times 24$ $48\times 48$ $96\times 96$ $192\times 192$
In the second experiment, we compare the preconditioned matrix $L^{-1}\tilde{S}_{p}$ to the non-preconditioned primal Schur complement $\tilde{S_{p}}$, in terms of the estimated 1-norm condition number. The results in Fig. 2(a) show that the condition number is significantly improved by the preconditioning. We note that the 1-norm condition number is significantly higher than the standard 2-norm condition number, yet easier to estimate numerically. We use the Matlab function condest for this estimation.
In the third experiment we compare the GMRES iteration count required for the solution of the original system $\tilde{S}_{p}p^{\prime}={\bf b}$ and the preconditioned system $L^{-1}\tilde{S}_{p}p^{\prime}=L^{-1}{\bf b}$. Note that during a full solution of the problem, the right hand side ${\bf b}$ changes at any time step. Here, we solved the systems for a random right hand side to demonstrate the improvement of the conditioning in terms of iteration count. The results shown in Fig. 2(b) demonstrate that the preconditioned system can be solved in just a few iterations regardless of the grid size, while the original system requires hundreds of iterations for very small grids and thousands for larger grids. This scalability property enables performing the simulations on dense grids, as shown in the next subsection.
(a) Conditioning
(b) Iterations
Figure 2: The 1-norm condition number per number of cells of the original and preconditioned matrices for the pressure corrections, and GMRES iteration count required for the solution of a system with a random right-hand-side for the corresponding matrices.
5.2 Three-dimensional experiments: an oscillating solid sphere
An oscillating solid sphere of diameter $D$ placed within a $4D\times 4D\times 6D$ box filled with a quiescent fluid (see Fig. 3(a)) is considered. The sphere oscillates vertically with velocity and position given by:
@@ 5.2 Three-dimensional experiments: an oscillating solid sphere
To demonstrate the efficiency of the preconditioner, we measure the average number of BiCGStab iterations required for convergence of the preconditioned system to a $10^{-12}$ tolerance, for different Reynolds numbers and grid sizes, as shown in Table 2. The average is taken over three different $A/D$ ratios: 0.5, 1.0, and 1.5, and over four distinct locations of the sphere during its periodic oscillation, i.e., at the lowest point of the stroke (bottom dead center), at the highest point of the stroke (top dead center), and at the two midpoints of the stroke as the sphere moves upwards and downwards. We observe that the iteration count remains nearly constant and remarkably low, regardless of the Reynolds number, grid size, $A/D$ ratio, and location of the sphere during its oscillatory motion.
Iteration count for a 3D oscillating sphere
Grid size $Re=10$ $Re=20$ $Re=50$ $Re=100$ $Re=200$ $Re=300$
$50\times 50\times 75$ 4.0 4.0 4.0 4.0 4.0 4.0
$100\times 100\times 150$ 4.25 4.0 4.0 4.17 4.17 4.0
$200\times 200\times 300$ 4.33 4.25 4.17 4.33 4.33 4.67
$400\times 400\times 600$ 4.17 4.0 4.08 4.08 4.33 4.25
Table 2: Average BiCGStab iteration number for the convergence of the preconditioned system to a $10^{-12}$ tolerance, averaged over different sphere locations and $A/D$ ratios.
(a) Time scalability
(b) Memory scalability
Figure 5: Wall-clock time per time-step and memory consumption for different grid sizes. The results were averaged over three $A/D$ ratio values: $0.5,1.0$ and $1.5$, and over six Reynolds number values: $10,20,50,100,200$ and $300$. The solid line is the linear regression line of the measurement points.
Finally, we measure the overall scalability properties of the algorithm. All the experiments were performed on a standard Linux server with 64GB DDR3 shared memory and 2 Intel Xeon 12C processors, 24 threads each (48 threads in total). Fig. 5 presents the time and memory consumption, averaged over measurements for different Reynolds numbers and different $A/D$ ratios. A regression line was added to analyze the dependence of time and memory consumption on the grid size. Our results show that the computational time (for a single time step) is of order $O(N^{1.1})$, where $N$ is the number of cells. The results are consistent with the complexity of the direct solver from [19], as estimated in [20], i.e., the direct solver takes $O(N^{1.33})$ in 3D, which is further decreased by using the Thomas algorithm in one of the dimensions. The results also demonstrate that the memory consumption is sub-linear. Moreover, for the largest grid examined—consisting of about 384 million unknowns—the method required only 25 GB of memory, an amount that can be found on a standard workstation and even a powerful laptop. For the sake of comparison, when the same problem was solved by [15], the memory consumption exceeded 128 GB for low $Re$ values. The performance of our algorithm does not depend on either time step or $Re$ values, as long as the CFL number is sufficiently small. [^4] We used a CFL number that does not exceed 0.1.
5.3 Three-dimensional experiments: an oscillating porous sphere
In the previous subsections we have established the scalability properties of our preconditioner for the case of a single body. In this subsection we preform additional experiments to further demonstrate the capabilities of the developed method for simulation of multiple moving bodies. In the view of Remark 4.1, the number of immersed bodies might affect the sparsity of $B$, and thus the scattering of the eigenvalues of the preconditioned system. The aim of this subsection is to demonstrate that the same scalable behaviour holds, regardless of the number of bodies.
@@ 5.3 Three-dimensional experiments: an oscillating porous sphere
The details regarding the radii and packing arrangements for each configuration of sub-spheres are given in Table 3. The table also includes the corresponding porosity values $\varepsilon$, defined as the ratio of void volume to total volume ($\varepsilon=V_{void}/V_{total}$). For numerical stability in the IBM implementation, each sub-sphere’s radius was reduced by half a grid spacing to ensure proper surface contact resolution while avoiding ill-conditioning. The physical configurations of the bounding spheres, containing 7 and 14 sub-spheres, are illustrated in Figs. 6(a) and 6(b), respectively. A spherical wire-frame is shown in these figures to indicate the boundary of the theoretical bounding sphere.
@@
The analysis of Figs. 7(a), 7(c) and 7(e) demonstrates that sub-spheres experience different temporal forces, with their magnitudes and patterns determined by vertical position within the bounding sphere. Sub-spheres at equal vertical positions experience identical temporal forces regardless of their horizontal coordinates, confirming the axial nature of the force distribution. The temporal evolution of forces on individual sub-spheres exhibits clear asymmetry, quantified by differences between positive and negative force peaks. Sub-spheres #1, #2, and #3, positioned below the bounding sphere’s center, experience negative forces approximately 50% larger than their positive counterparts during downward motion, with reduced forces during upward motion due to shielding from upper sub-spheres. Sub-sphere #7, positioned at the top edge of the bounding sphere, displays an inverse asymmetry pattern, with positive force peaks exceeding negative values by approximately 40%. Sub-spheres #4, #5, and #6, located slightly above the center and attached to the outer boundary of the bounding sphere, remain unshielded by other sub-spheres throughout the oscillation cycle. This unshielded position results in nearly symmetric forces about zero, with peak-to-peak variations not exceeding 10%. The magnitude of these forces systematically decreases with increasing Reynolds number across all cases, while maintaining consistent phase relationships relative to the bounding sphere’s position. This reduction in force amplitude demonstrates the diminishing effect of viscous forces at higher Reynolds numbers, though the fundamental force distribution patterns persist.
@@
The behavior of external forces in the case of 14 sub-spheres exhibits similar patterns while revealing additional features due to the more complex internal structure. The analysis of Figs. 7(b), 7(d) and 7(f) confirms that sub-spheres at equal vertical positions experience identical temporal forces regardless of their horizontal coordinates. Sub-spheres #1 through #4, positioned at $z\approx-0.249$ below the bounding sphere’s center, experience negative forces approximately 45% larger than their positive counterparts during downward motion, with reduced forces during upward motion due to shielding from upper sub-spheres. Sub-spheres #5 through #8, located near the equatorial plane at $z\approx-0.030$, show intermediate behavior with force asymmetry less pronounced than for the bottom spheres. Their position, slightly below the geometric center, results in force patterns that reflect partial shielding effects from both upper and lower neighboring spheres. The force distribution of sphere #9, uniquely positioned at the exact geometric center ($z=0$) of the confining sphere, demonstrates nearly symmetric behavior about zero, with peak-to-peak variations not exceeding 5%, due to balanced shielding from surrounding sub-spheres. Sub-spheres #10 through #13, located at $z\approx 0.184$, display moderate force asymmetry with positive peaks approximately 30% larger than negative ones. Sub-sphere #14, positioned at $z\approx 0.338$ at the top edge of the bounding sphere, exhibits the strongest asymmetry in the upper region, with positive force peaks exceeding negative values by approximately 35%. As in the previous case, these asymmetric force patterns persist across all examined $Re$ values, with diminishing magnitudes at higher Reynolds numbers.
@@
Looking at Fig. 8, we observe three key trends in the drag coefficient evolution. First, the magnitude of $C_{D}$ systematically decreases with increasing Reynolds numbers, for both 7 and 14 sub-sphere arrays. This trend directly reflects diminishing viscous effects at higher $Re$ values, resulting in reduced drag forces. Second, the 14 sub-sphere configuration exhibits consistently higher $C_{D}$ values compared to the 7 sub-sphere array across all Reynolds numbers, approaching the drag coefficient values that characterize a non-porous sphere, as reported in our previous work (see [15], Fig. 7). This behavior is expected, as increasing the number of sub-spheres reduces the array’s porosity, making it more similar to a solid sphere. Third, a persistent phase lag exists between the sphere position and drag coefficient evolution. This phase lag arises from the fundamental relationship between force and motion in oscillatory flows: the drag force is primarily dependent on the velocity rather than the displacement. At low $Re$ values, the force reaches its maximum when the velocity is highest (at zero displacement) and its minimum when the velocity is zero (at maximum displacement), creating an inherent quarter-period phase shift. This kinematic relationship persists across all $Re$ values, with additional inertial effects becoming more pronounced at higher Reynolds numbers but not altering the basic phase relationship. Note that a similar phase lag is present in Fig. 4 from the previous subsection, corresponding to a single oscillating sphere.
@@
To gain deeper insight into the fluid dynamics associated with the transverse motion of porous spheres, we visualize the vortical structures generated by the flow. The visualization utilizes isosurfaces of the $Q$-criterion ($Q=0.05$), where $Q$ represents the second invariant of the velocity gradient tensor. This value of $Q$ was chosen to effectively capture coherent vortical structures while filtering out weaker vorticity regions, thereby highlighting the dominant flow features. The evolution of vortical structures over one oscillation period is shown in Figs. 9 and 10 for configurations containing 7 and 14 sub-spheres, respectively. In both configurations, a clear Reynolds number dependence is observed in the persistence of vortical structures. At $Re=100$, vortices dissipate rapidly due to dominant viscous diffusion. As $Re=300$, vortical structures maintain their coherence over longer periods and travel further from the sphere surface before dissipating, reflecting the diminishing role of viscous effects relative to inertial forces.

### v2

@@ 5 Theory: a spectral bound on the preconditioned system
In this section, we analyze the use of the Laplacian $L$ as a preconditioner for the Schur complement $S_{p}$ of the coupled pressure–force-density system (20) or (27). A good preconditioner should resemble the original matrix, in the sense of spectral equivalence, i.e., the spectrum of the preconditioned system should be bounded. Theorem 5.1 shows this property.
@@
has only solutions that satisfy $\lambda\in[0,1]$. Observe that
$B^{T}C^{-1}B=G^{T}R(R^{T}R)^{-1}R^{T}G=G^{T}PG$
where $P\coloneqq R(R^{T}R)^{-1}R^{T}$ is an orthogonal projector (namely, $P^{2}=P$ and $P^{T}=P$). To bound $\lambda$, we bound the generalized Rayleigh quotient
$\frac{\left\langle B^{T}C^{-1}B{\bf v},{\bf v}\right\rangle}{\left\langle L{\bf v},{\bf v}\right\rangle}=\frac{\left\langle G^{T}PG{\bf v},{\bf v}\right\rangle}{\left\langle G^{T}G{\bf v},{\bf v}\right\rangle}=\frac{\left\langle PG{\bf v},G{\bf v}\right\rangle}{\left\langle G{\bf v},G{\bf v}\right\rangle},\quad{\bf v}\in\mathbb{R}^{m}.$ (39)
Since it is an orthogonal projector, the only eigenvalues of $P$ are 0 and 1, and $\text{Null}(P)=\text{Range}(P)^{\perp},$ see, e.g., [46]. Particularly, every vector can be uniquely decomposed into a sum of orthogonal vectors in the nullspace and in the range of $P$, e.g., we can write $G{\bf v}={\bf v}_{1}+{\bf v}_{2}$, where ${\bf v}_{1}\in\text{Null}(P)$, ${\bf v}_{2}\in\text{Range}(P)$ and $\left\langle{\bf v}_{1},{\bf v}_{2}\right\rangle=0.$ Consequently,
@@
$0\leq\frac{\left\langle PG{\bf v},G{\bf v}\right\rangle}{\left\langle G{\bf v},G{\bf v}\right\rangle}=\frac{\|{\bf v}_{2}\|^{2}}{\|{\bf v}_{1}+{\bf v}_{2}\|^{2}}=\left(\frac{\|P({\bf v}_{1}+{\bf v}_{2})\|}{\|{\bf v}_{1}+{\bf v}_{2}\|}\right)^{2}\leq\|P\|^{2}_{2}=\left(\rho(P)\right)^{2}=1$ (41)
@@
The assumption of symmetry is not particularly restrictive, as it is natural to take the same discrete Dirac delta functions for the regularization and interpolation, and the same discretization for the gradient and divergence. In practice, some implementation of the possible BC might lead to a non-symmetric Laplacian. Nevertheless, even when the Laplacian slightly differs from $G^{T}G$, the spectral result of Theorem 5.1 seems to hold in practice, as demonstrated in the next section.
@@
As an immediate corollary of Theorem 5.1, the error iteration matrix $T=I_{m}-L^{-1}S_{p}$ of the preconditioned system satisfies $\rho(T)\leq 1$. When this inequality is strong, convergence is guaranteed. A Krylov iterative method typically further accelerates the convergence and might converge in just a few iterations even when the spectral radius of the error iteration matrix is even greater than 1. However, the convergence rate of a Krylov iterative method is determined by the scattering of the eigenvalues, which can depend on the specific configuration. In typical cases, the number of Lagrangian points is very small compared to the number of Eulerian cells. In such a case,
@@ 6 Results
In this section we demonstrate the efficiency of our method in solving various problems. In Subsection 6.1 we present several model problems in 3D, and verify our method by comparing the simulation results to existing references. In subsection 6.2 we demonstrate the computational efficiency of our preconditioning approach in 2D and in 3D.
6.1 Verification study
6.1.1 Model problem 1: a transversely oscillating solid sphere (one-way coupling)
A transversely oscillating solid sphere of diameter $D$ placed within a $4D\times 4D\times 6D$ box filled with a quiescent fluid (see Fig. 2(a)) is considered. The sphere oscillates vertically with velocity and position given by:
@@ 6.1.1 Model problem 1: a transversely oscillating solid sphere (one-way coupling)
6.1.2 Model problem 2: multiple packed spheres (one-way coupling)
In the previous subsection we verified our method for the case of a single body. In this subsection we preform additional experiments to further demonstrate the capabilities of the developed method for simulation of multiple moving bodies.
@@ 6.1.2 Model problem 2: multiple packed spheres (one-way coupling)
The details regarding the radii and packing arrangements for each configuration of sub-spheres are given in Table 1. The table also includes the corresponding porosity values $\varepsilon$, defined as the ratio of void volume to total volume ($\varepsilon=V_{void}/V_{total}$). For numerical stability in the IBM implementation, each sub-sphere’s radius was reduced by half a grid spacing to ensure proper surface contact resolution while avoiding ill-conditioning. The physical configurations of the bounding spheres, containing 7 and 14 sub-spheres, are illustrated in Figs. 4(a) and 4(b), respectively. A spherical wire-frame is shown in these figures to indicate the boundary of the theoretical bounding sphere.
@@
The analysis of Figs. 5(a), 5(c) and 5(e) demonstrates that sub-spheres experience different temporal forces, with their magnitudes and patterns determined by vertical position within the bounding sphere. Sub-spheres at equal vertical positions experience identical temporal forces regardless of their horizontal coordinates, confirming the axial nature of the force distribution. The temporal evolution of forces on individual sub-spheres exhibits clear asymmetry, quantified by differences between positive and negative force peaks. Sub-spheres #1, #2, and #3, positioned below the bounding sphere’s center, experience negative forces approximately 50% larger than their positive counterparts during downward motion, with reduced forces during upward motion due to shielding from upper sub-spheres. Sub-sphere #7, positioned at the top edge of the bounding sphere, displays an inverse asymmetry pattern, with positive force peaks exceeding negative values by approximately 40%. Sub-spheres #4, #5, and #6, located slightly above the center and attached to the outer boundary of the bounding sphere, remain unshielded by other sub-spheres throughout the oscillation cycle. This unshielded position results in nearly symmetric forces about zero, with peak-to-peak variations not exceeding 10%. The magnitude of these forces systematically decreases with increasing Reynolds number across all cases, while maintaining consistent phase relationships relative to the bounding sphere’s position. This reduction in force amplitude demonstrates the diminishing effect of viscous forces at higher Reynolds numbers, though the fundamental force distribution patterns persist.
@@
The behavior of external forces in the case of 14 sub-spheres exhibits similar patterns while revealing additional features due to the more complex internal structure. The analysis of Figs. 5(b), 5(d) and 5(f) confirms that sub-spheres at equal vertical positions experience identical temporal forces regardless of their horizontal coordinates. Sub-spheres #1 through #4, positioned at $z\approx-0.249$ below the bounding sphere’s center, experience negative forces approximately 45% larger than their positive counterparts during downward motion, with reduced forces during upward motion due to shielding from upper sub-spheres. Sub-spheres #5 through #8, located near the equatorial plane at $z\approx-0.030$, show intermediate behavior with force asymmetry less pronounced than for the bottom spheres. Their position, slightly below the geometric center, results in force patterns that reflect partial shielding effects from both upper and lower neighboring spheres. The force distribution of sphere #9, uniquely positioned at the exact geometric center ($z=0$) of the confining sphere, demonstrates nearly symmetric behavior about zero, with peak-to-peak variations not exceeding 5%, due to balanced shielding from surrounding sub-spheres. Sub-spheres #10 through #13, located at $z\approx 0.184$, display moderate force asymmetry with positive peaks approximately 30% larger than negative ones. Sub-sphere #14, positioned at $z\approx 0.338$ at the top edge of the bounding sphere, exhibits the strongest asymmetry in the upper region, with positive force peaks exceeding negative values by approximately 35%. As in the previous case, these asymmetric force patterns persist across all examined $Re$ values, with diminishing magnitudes at higher Reynolds numbers.
@@
Looking at Fig. 6, we observe three key trends in the drag coefficient evolution. First, the magnitude of $C_{D}$ systematically decreases with increasing Reynolds numbers, for both 7 and 14 sub-sphere arrays. This trend directly reflects diminishing viscous effects at higher $Re$ values, resulting in reduced drag forces. Second, the 14 sub-sphere configuration exhibits consistently higher $C_{D}$ values compared to the 7 sub-sphere array across all Reynolds numbers, approaching the drag coefficient values that characterize a non-porous sphere, as reported in our previous work (see [47], Fig. 7). This behavior is expected, as increasing the number of sub-spheres reduces the array’s porosity, making it more similar to a solid sphere. Third, a persistent phase lag exists between the sphere position and drag coefficient evolution. This phase-lag arises from the fundamental relationship between force and motion in oscillatory flows: the drag force is primarily dependent on the velocity rather than the displacement. At low $Re$ values, the force reaches its maximum when the velocity is highest (at zero displacement) and its minimum when the velocity is zero (at maximum displacement), creating an inherent quarter-period phase shift. This kinematic relationship persists across all $Re$ values, with additional inertial effects becoming more pronounced at higher Reynolds numbers but not altering the basic phase relationship. Note that a similar phase-lag is present in Fig. 3 from the previous subsection, corresponding to a single oscillating sphere.
@@
To gain deeper insight into the fluid dynamics associated with the transverse motion of porous spheres, we visualize the vortical structures generated by the flow. The visualization utilizes isosurfaces of the $Q$-criterion ($Q=0.05$), where $Q$ represents the second invariant of the velocity gradient tensor. This value of $Q$ was chosen to effectively capture coherent vortical structures while filtering out weaker vorticity regions, thereby highlighting the dominant flow features. The evolution of vortical structures over one oscillation period is shown in Figs. 7 and 8 for configurations containing 7 and 14 sub-spheres, respectively. In both configurations, a clear Reynolds number dependence is observed in the persistence of vortical structures. At $Re=100$, vortices dissipate rapidly due to dominant viscous diffusion. As $Re=300$, vortical structures maintain their coherence over longer periods and travel further from the sphere surface before dissipating, reflecting the diminishing role of viscous effects relative to inertial forces.